OpenStudy (moss65):

joint probability density function f(x,y)=6(1-y) for 0<=x<=y<=1, and =0 elsewhere. Find P(y<=1/2|x<=3/4).

OpenStudy (anonymous):

I believe it would just be: \[ \int_{-\infty}^{1/2}\int_{-\infty}^{3/4}f(x,y)~dx~dy \]

OpenStudy (anonymous):

You can change the lower bounds to \(0\) since \(\int_{-\infty}^0\) will be \(0\).

OpenStudy (anonymous):

Wait hold on. this is conditional probability?

OpenStudy (moss65):

Yes, it is

OpenStudy (anonymous):

Well, what I do gives the probability that both are under their respective value.

OpenStudy (anonymous):

So we need to use: \[ \Pr(Y|X) = \frac{\Pr(Y\cap X)}{\Pr(X)} \]

OpenStudy (anonymous):

We have the \(Y\cap X\) part down.

OpenStudy (moss65):

yes, but what next?

OpenStudy (anonymous):

We need to find \(\Pr(X)\). I believe it would be given by \(f_X(x<3/4)\).

OpenStudy (anonymous):

Do you know how to find that?

OpenStudy (moss65):

f_x(x)=3(1-x)^2

OpenStudy (anonymous):

Do you sort of understand where I am getting these things?

OpenStudy (anonymous):

I think you might have written \(f(x,y)\) wrong, because it has no \(x\).

OpenStudy (moss65):

not quite. I wrote it correctly. before this question, i had to find the marginal density functions for X and Y. they are f(x)=3(1-x)^2, f(y)=6y(1-y). now I have to find the conditional probability shown above. it is my "weak place" Textbook examples age giving examples with X or Y given, sort of P(X<=1|Y=2)

OpenStudy (anonymous):

Hmm, now that I think about it, the \(x\leq y\) part does make things a bit tricky.

OpenStudy (moss65):

i think so

OpenStudy (anonymous):

I guess you had to use \(f_X (x)= \int_x^1f(x,y)~dy\) And then \(f_Y(y) = \int_0^y f(x,y)~dx\)

OpenStudy (moss65):

I used it to find f(x) and f(y). Now i need to use somehow f(x,y)/f(x)

OpenStudy (anonymous):

If we assume that \(x\leq 3/4\)...

OpenStudy (anonymous):

actually, that really doesn't help us much at all, lol.

OpenStudy (anonymous):

But I'm not quite sure that is correct either.

OpenStudy (anonymous):

Okay, now I'm thinking:\[ \int_0^{3/4}\int_x^{1/2}f(x,y)~dy~dx \]

OpenStudy (anonymous):

Hmmm, I think that ought to keep us so that \(x\leq y\)... don't you?

OpenStudy (anonymous):

But maybe not, since we are letting \(x\) to up to \(3/4\)...

OpenStudy (moss65):

I don't know. I know the answer (32/63) but i have no idea how to get it.

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay, I have another idea... what if we have \(y = t+x\)

OpenStudy (moss65):

and?

OpenStudy (anonymous):

Then \(0\leq t\leq1-x\)

OpenStudy (anonymous):

I am just trying to see if we can change this dependency on \(x\).

OpenStudy (anonymous):

So we change: \[ \int_0^{3/4}\int_x^{1/2}f(x,y)~dy~dx =\int_0^{3/4}\int_0^{1/2-x} f(x,t+x)~dt~dx \]

OpenStudy (moss65):

for me, it still looks like P(x,y) or P(x,t)

OpenStudy (moss65):